OpenStudy (anonymous):

if sqrt 3 is a zero of the polynomial g(x)= 2x^2- ax, what must the value of a equal if sqrt 3 is a zero of the polynomial g(x)= 2x^2- ax, what must the value of a equal @Mathematics

myininaya (myininaya):

\[g(\sqrt{3})=0\]

myininaya (myininaya):

\[g(\sqrt{3})=2(\sqrt{3})^2-a \sqrt{3}=0\]

myininaya (myininaya):

\[2(3)- a \sqrt{3}=0\] i will let you do the rest

OpenStudy (anonymous):

haha wow i really over thought it. thanks

myininaya (myininaya):

np